OpenStudy (anonymous):

lim 2 - (1/2)^x x--->3

OpenStudy (vishweshshrimali5):

Just put x = 3 in the expression : 2-(1/2)^x.

OpenStudy (anonymous):

my bad it's as x approaches infinity

OpenStudy (vishweshshrimali5):

Ohh

OpenStudy (vishweshshrimali5):

Now then you will have to use a particular formula. Let me recall it.

OpenStudy (anonymous):

no its using a graphing calculator.

OpenStudy (vishweshshrimali5):

Ohh sorry I do not have much idea about it.

OpenStudy (anonymous):

as x approaches infinity, (1/2)^x approaches 0, whole limit approaches 2

OpenStudy (anonymous):

1/2=0.5 1/2^2=0.25 1/2^3=0.125... as n tends to infinity 1/2^n tends to 0